# Conditional PED-ANOVA: Hyperparameter Importance in Hierarchical & Dynamic Search Spaces

### v1

@@ 6. Conclusion
We studied HPI estimation in conditional search spaces, where hyperparameters may appear, disappear, or change domain depending on other choices. We identified a fundamental pitfall of standard local HPI: conditioning effects can leak into conditioned hyperparameters, leading to misleading importances. To address this, we defined conditional local HPI and proposed condPED-ANOVA, a closed-form, PED-ANOVA-style estimator that correctly handles conditional search spaces while retaining efficiency. Empirically, condPED-ANOVA yielded meaningful importances consistent with the underlying conditional structure, whereas naive extensions of existing estimators, including PED-ANOVA, often violated conditional semantics and produced distorted importances. We believe condPED-ANOVA serves as a practical and principled foundation for reliable HPI analysis in HPO workflows with conditional search spaces.
@@ A.1. Contrastive Hyperparameter Importance
Beyond global variance-based rankings, several lines of work emphasize selection, contrastive explanations, or local sensitivity. Hutter et al. (2013) study forward selection to identify a small subset of influential inputs (parameters and instance features) by incrementally adding dimensions that most improve surrogate predictive accuracy, followed by a drop-one analysis to rank selected inputs. In parallel, ablation analysis (Fawcett and Hoos, 2016) explains performance differences between two high-quality configurations by constructing a path that sequentially ablates parameter changes and quantifies their contributions, yielding a contrastive, configuration-to-configuration notion of importance. Biedenkapp et al. (2017) accelerate this paradigm via ablation with surrogates, replacing expensive evaluations with model-based predictions to achieve large speedups while maintaining explanatory utility. To further focus on neighborhood behavior, CAVE (Biedenkapp et al., 2019) introduces local parameter importance, which measures variance in predicted performance when varying one parameter around a specific configuration, and empirically argues that local and global importance can disagree and thus complement each other.
@@ A.2. Importance Across Datasets and Tunability
Another line of work studies importance beyond a single dataset or run. Van Rijn and Hutter (2018) aggregates evidence across many datasets (via OpenML metadata) to identify which hyperparameters tend to matter most for an algorithm in general, and to infer priors over promising values; this reframes HPI as a meta-learning problem and supports more data-driven space design and warm-starting. Complementarily, (Probst et al., 2019) introduce tunability measures that quantify how much performance can improve by tuning, providing a practitioner-facing notion of “which hyperparameters are worth tuning” and offering data-based defaults and empirical assessments across datasets and algorithms.
@@ B.1. Extended One-Dimensional Domain
$\mathsf{S}^{(d)}\coloneqq\bigsqcup_{i=1}^{\;\;\;K^{(d)}}\left(\{i\}\times\mathcal{Z}_{i}^{(d)}\right),$
@@ B.2. Rigorous Formulation of condPED-ANOVA
$\alpha_{i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(A_{i}^{(d)}),\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(A_{i}^{(d)}).$
@@
$\mu_{\gamma^{\prime},\,i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(\cdot\mid A_{i}^{(d)}),\quad\text{and}\quad\mu_{\gamma,\,i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(\cdot\mid A_{i}^{(d)}).$
@@
With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
@@
$D_{\mathrm{PE}}(\nu\|\mu)\coloneqq\int\left(\frac{d\nu}{d\mu}-1\right)^{\!2}\,d\mu,$
for probability measures $\nu$ and $\mu$ with $\nu\ll\mu$. Note that the summation in Equation 14 is restricted to regimes with $\beta_{i}^{(d)}>0$; regimes absent from the top-$\gamma$ set contribute neither to the within-regime variance nor to the divergence. If $A_{i}^{(d)}$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma,\,i}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}$ are both degenerate on $A_{i}^{(d)}$, implying $D_{\mathrm{PE}}(\mu_{\gamma^{\prime},\,i}^{(d)}\|\mu_{\gamma,\,i}^{(d)})=0$. When $K^{(d)}=1$, we have $\alpha_{1}=\beta_{1}=1$ and Equation 14 reduces to the original PED-ANOVA expression.
@@
$\mathbb{P}(S\in B,\,Y=1)$ $=\mathbb{P}(Y=1)\,\mathbb{P}(S\in B\mid Y=1)$
@@
$\mathbb{P}(Y=1\mid S=s)\;=\;\kappa\,\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)$
@@
$v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I}\!\left[\mathrm{Var}\!\left(\mathbb{P}(Y=1\mid S)\,\middle|\,I\right)\right].$
@@
$v_{\gamma,\mathrm{within}}^{(d)}$ $=\sum_{i:\,\beta_{i}^{(d)}>0}\beta_{i}^{(d)}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\kappa\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
$=\kappa^{2}\sum_{i:\,\beta_{i}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right).$
@@
$\mathbb{E}_{s\sim\mu_{\gamma,i}^{(d)}}\left[\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right]=\int\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}\,d\mu_{\gamma,\,i}^{(d)}=\mu_{\gamma^{\prime},\,i}^{(d)}(\mathsf{S}^{(d)})=1.$
@@
Appendix C Additional Results and Discussion on Numerical Experiments
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 5. condPED-ANOVA HPI computed for the synthetic objectives. For the objective with conditional activation (Equation 6), the gating hyperparameter $c$ determines which branch is active: $x$ is present only when $c<0.5$, whereas $y$ is present only when $c\geq 0.5$. For the objective with regime-dependent domains (Equation 9), the gating hyperparameter $c$ determines the domain of the hyperparameters $x$ and $y$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
C.1. condPED-ANOVA Results on Synthetic Objectives with Different $\gamma$ Values
The results of condPED-ANOVA for the objective with conditional activation (Equation 6) and the objective with regime-dependent domains (Equation 9), with $\gamma$ set to $0.75$ and $0.5$, are shown in Figure 5.
The results share some similarities with the higher-$\gamma^{\prime}$ range in Figure 1, which reflects the fact that varying $\gamma$ changes the region over which HPI is computed. However, the behavior is not a simple zoom-in because the definition of the top-$\gamma$ region itself changes. In particular, with $\gamma=0.5$ in the disjoint-domain setting (Figure 5, right panel), once we restrict attention to the better half of the evaluations (i.e., the region where $c<0.5$ is already selected), adjusting $c$ no longer changes the objective value. Accordingly, the importance of $c$ becomes small in this regime, similar to the importance of $y$.
C.2. Comparison with Existing Methods
Figure 6. Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation (Equation 6) under the overlapping domain setting (Equation 8). The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
The results for the baseline methods described in Section 5.4 for the conditional-activation objective with the overlapping-domain setting (Equations 6 and 8) and the regime-dependent domain setting (Equations 9 and 10) are shown in Figures 6 and 7, respectively.
First, for the conditional-activation objective, we observe the same trend as in the disjoint-domain setting (Equations 6 and 7). PED-ANOVA with filtering fails to distinguish between $x$ and $y$, while PED-ANOVA with imputation assigns $y$ an importance comparable to $c$ in the small-$\gamma^{\prime}$ regime, where $y$ should be inactive and thus cannot improve performance. Similarly, f-ANOVA with filtering and MDI with filtering again yield nearly identical HPI values for all hyperparameters, and the remaining baselines also tend to assign almost the same importance to $x$ and $y$.
Next, for the regime-dependent domain setting, PED-ANOVA with expansion also behaves undesirably: although $c$ determines whether a configuration falls into the top half, the importance of $c$ around $\gamma^{\prime}\approx 0.5$ becomes smaller than that of $x$ and $y$. This failure is caused by expansion, which collapses the regime-specific ranges of $x$ and $y$ into a single unified domain and thereby obscures the effect of range switching, preventing accurate estimation of $x$ and $y$ in the regimes where their domains change. Moreover, for f-ANOVA, MDI, and SHAP with expansion, $y$ is assigned larger importance than $x$, even though $x$ should contribute more to the objective. This is because forcibly expanding the domain distorts the relationship between the variables and the objective across regimes, leading to misleading HPI estimates.
These results indicate that the naive extensions do not provide meaningful HPI estimates that capture the difference between the contributions of $x$ and $y$ in this conditional setting, not only for the conditional-activation objective under the disjoint-domain setting (as shown in Section 5.4.2), but also under the overlapping-domain setting and in the regime-dependent domain setting.
C.3. Runtime Comparison
We compare the runtime of condPED-ANOVA against baseline HPI methods equipped with naive extensions on the synthetic objective with regime-dependent domains (Equation 9 and Equation 10). For each sample size from $2^{9}$ to $2^{17}$, we measure the wall-clock time required to compute HPI, and report the results in Figure 8. f-ANOVA is reported only up to $2^{11}$, since it could not be executed at larger sample sizes due to excessive memory consumption. Due to its substantial memory consumption, f-ANOVA could not be executed for larger sample sizes and is reported only up to $2^{11}$. Similarly, SHAP is reported only up to $2^{15}$ because it was prohibitively slow at larger sample sizes.
(c) Regime-dependent domains (Equation 9 and Equation 10)
C.4. Ablation Study on the Effect of Regime-Level Weighting
The ablation study results obtained by using a naive aggregation scheme (Equation 12) instead of the closed-form expression derived in Theorem 4.2 are shown in Figure 4. For reference, we also show the HPI values before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$, in Figure 9.
Appendix D Detailed Experimental Setup
D.1. Details for Common Experimental Setup
@@
D.2. Implementation Details for Baselines
@@
Algorithm 2 Naive Extension: Filtering (Conditional Presence)
@@
8:   $\text{HPI}\leftarrow d\text{-th value of }\textsc{BaselineHPI}(\mathcal{D}_{\mathrm{act}}^{(d)})$
@@
Algorithm 3 Naive Extension: Imputation (Conditional Presence)
@@
Algorithm 4 Naive Extension: Expansion (Regime-Specific Ranges)

### v2

@@ 7. Conclusion
We studied HPI estimation in conditional search spaces, where hyperparameter activation and domains depend on others. We identified a key pitfall of standard local HPI: conditioning effects can leak into conditioned hyperparameters, leading to misleading HPIs. To address this, we introduced conditional local HPI and proposed condPED-ANOVA, a closed-form, PED-ANOVA-style estimator that correctly handles conditionality while retaining efficiency. Experiments showed that condPED-ANOVA produces meaningful HPIs under conditional search spaces, whereas naive extensions of existing estimators often yield misleading HPIs. We believe condPED-ANOVA serves as a practical and principled foundation for reliable HPI analysis in HPO workflows with conditional search spaces.
@@ A.1. Contrastive Hyperparameter Importance
Beyond global variance-based rankings, several lines of work emphasize selection, contrastive explanations, or local sensitivity. Hutter et al. (2013) study forward selection to identify a small subset of influential inputs (parameters and instance features) by incrementally adding dimensions that most improve surrogate predictive accuracy, followed by a drop-one analysis to rank selected inputs. In parallel, ablation analysis (Fawcett and Hoos, 2016) explains performance differences between two high-quality configurations by constructing a path that sequentially ablates parameter changes and quantifies their contributions, yielding a contrastive, configuration-to-configuration notion of importance. Biedenkapp et al. (2017) accelerate this paradigm via ablation with surrogates, replacing expensive evaluations with model-based predictions to achieve large speedups while maintaining explanatory utility. To further focus on neighborhood behavior, CAVE (Biedenkapp et al., 2019) introduces local parameter importance, which measures variance in predicted performance when varying one parameter around a specific configuration, and empirically argues that local and global importance can disagree and complement each other.
@@ A.2. Importance Across Datasets and Tunability
Another line of work studies the importance beyond a single dataset or run. van Rijn and Hutter (2018) aggregates evidence across many datasets (via OpenML metadata) to identify which hyperparameters tend to matter most for an algorithm in general, and to infer priors over promising values; this reframes HPI as a meta-learning problem and supports more data-driven space design and warm-starting. Complementarily, (Probst et al., 2019) introduce tunability measures that quantify how much performance can improve by tuning, providing a practitioner-facing notion of “which hyperparameters are worth tuning” and offering data-based defaults and empirical assessments across datasets and algorithms.
@@ B.1. Extended One-Dimensional Domain
(19) $\mathsf{S}^{(d)}\coloneqq\bigsqcup_{i=1}^{\;\;\;K^{(d)}}\left(\{i\}\times\mathcal{Z}_{i}^{(d)}\right),$
@@ B.2. Rigorous Formulation of condPED-ANOVA
(20) $\alpha_{i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(A_{i}^{(d)}),\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(A_{i}^{(d)}).$
@@
(21) $\mu_{\gamma^{\prime},\,i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(\cdot\mid A_{i}^{(d)}),\quad\text{and}\quad\mu_{\gamma,\,i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(\cdot\mid A_{i}^{(d)}).$
@@
With these definitions, we obtain the following closed-form estimation for the within-regime local marginal variance.
@@
(24) $D_{\mathrm{PE}}(\nu\|\mu)\coloneqq\int\left(\frac{d\nu}{d\mu}-1\right)^{\!2}\,d\mu,$
for probability measures $\nu$ and $\mu$ with $\nu\ll\mu$. Note that the summation in Equation 23 is restricted to regimes with $\beta_{i}^{(d)}>0$; regimes absent from the top-$\gamma$ set contribute neither to the within-regime variance nor to the divergence. If $A_{i}^{(d)}$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma,\,i}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}$ are both degenerate on $A_{i}^{(d)}$, implying $D_{\mathrm{PE}}(\mu_{\gamma^{\prime},\,i}^{(d)}\|\mu_{\gamma,\,i}^{(d)})=0$. When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}=\beta_{1}=1$, so Equation 23 reduces to the original PED-ANOVA expression.
@@
(25) $\mathbb{P}(S\in B,\,Y=1)$ $=\mathbb{P}(Y=1)\,\mathbb{P}(S\in B\mid Y=1)$
@@
(29) $\mathbb{P}(Y=1\mid S=s)\;=\;\kappa\,\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)$
@@
(30) $v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I}\!\left[\mathrm{Var}\!\left(\mathbb{P}(Y=1\mid S)\,\middle|\,I\right)\right].$
@@
(31) $v_{\gamma,\mathrm{within}}^{(d)}$ $=\sum_{i:\,\beta_{i}^{(d)}>0}\beta_{i}^{(d)}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\kappa\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
$=\kappa^{2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
@@
(32) $\mathbb{E}_{s\sim\mu_{\gamma,i}^{(d)}}\left[\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right]=\int\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}\,d\mu_{\gamma,\,i}^{(d)}=\mu_{\gamma^{\prime},\,i}^{(d)}(\mathsf{S}^{(d)})=1.$
@@
Appendix C Detailed Experimental Setup
C.1. Details for Common Experimental Setup
@@
C.2. Implementation Details for Baselines
@@
Algorithm 2 Naive Extension: Filtering for Conditional Presence
@@
8:   $\text{HPI}^{(d)}\leftarrow d\text{-th value of }\textsc{BaselineHPI}(\mathcal{D}_{\mathrm{act}}^{(d)})$
@@
Algorithm 3 Naive Extension: Imputation for Conditional Presence
@@
Algorithm 4 Naive Extension: Expansion for Regime-Dependent Domains
@@
C.3. Detailed Setup for the Experiments on Real-World Benchmarks (Section 6)
We use YAHPO Gym (Pfisterer et al., 2022) as a surrogate-based benchmark suite for HPO, which provides fast evaluations for diverse search spaces and datasets. In YAHPO Gym, a scenario specifies a specific search space and a set of target metrics, while different instances correspond to different datasets/tasks within that scenario.
YAHPO Gym includes the following scenario families:
  • rbv2: Classical ML pipelines with mixed and often hierarchical/conditional hyperparameter spaces across many OpenML datasets, supporting multi-fidelity via training-size/fraction.
  • nb301: NAS-Bench-301 (Zela et al., 2022) based neural architecture search with a high-dimensional categorical space and epoch-based fidelity, targeting validation accuracy and runtime.
  • lcbench: Learning-curve benchmark over many OpenML (Vanschoren et al., 2014) tasks with epoch-based fidelity, exposing multiple validation/test performance metrics and training time.
  • iaml: Interpretable AutoML scenarios that combine predictive performance with resource-usage and interpretability objectives, with multi-fidelity via training-size fractions.
Among these, we use the rbv2_super scenario because it is single-objective and contains conditional search spaces. We follow the benchmark suites proposed with YAHPO Gym, where instances are selected based on surrogate faithfulness and diversity across scenarios; in the single-objective suite, rbv2_super includes six instances (which correspond to OpenML dataset IDs): 1053, 1457, 1063, 1479, 15, and 1468. We conduct experiments on all six instances. We present results for instance ID 1053 in the main text (Section 6) and report the remaining results in Appendix G.
The remaining experimental settings are the same as those in the synthetic experiments (Sections 5.1 and C.1).
Appendix D Additional Results and Discussion on Experiments on Synthetic Problem (Section 5)
In Section 5, we presented the main results and discussion of experiments on synthetic objectives. In this section, we provide additional results and discussion that complement those in Section 5.
D.1. Runtime Comparison
We compare the runtime of condPED-ANOVA against baseline HPI methods on the synthetic objective with regime-dependent domains (Equation 15 and Equation 16). For each sample size from $2^{9}$ to $2^{17}$, we measure the wall-clock time required to compute HPI, and report the results in Figure 6. f-ANOVA is reported only up to $2^{11}$, since it could not be executed at larger sample sizes due to excessive memory consumption. Similarly, SHAP is reported only up to $2^{15}$ because it was prohibitively slow at larger sample sizes.
(c) Regime-dependent domains (Equation 15 and Equation 16)
Figure 7. condPED-ANOVA HPI computed for the synthetic objectives with varying $\gamma$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
(c) Regime-dependent domains (Equation 15 and Equation 16)
Figure 8. condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives with varying NN. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
Figure 9. Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation under the overlapping domain setting (Equations 12 and 14). The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
D.2. condPED-ANOVA Results with Different $\gamma$
The results of condPED-ANOVA for the synthetic objectives, with $\gamma$ set to $0.75$ and $0.5$, are shown in Figure 7.
The results share some similarities with the higher-$\gamma^{\prime}$ range in Figure 1, which reflects the fact that varying $\gamma$ changes the region over which HPI is computed. However, the behavior is not a simple zoom-in because the definition of the top-$\gamma$ region itself changes. In particular, with $\gamma=0.5$ in the disjoint-domain setting (Figure 7, right panel), once we restrict attention to the better half of the evaluations (i.e., the region where $c<0.5$ is already selected), adjusting $c$ no longer changes the objective value. Accordingly, the HPI of $c$ becomes small in this regime, similar to the HPI of $y$.
D.3. condPED-ANOVA Results with Different NN
The results of condPED-ANOVA for the synthetic objectives, with $N$ set to $100$ and $500$, are shown in Figure 8.
From this, we observe that reducing $N$ slightly increases the standard error. However, the resulting HPI estimates remain reasonable and are largely consistent with those obtained when $N=1000$. These results suggest that condPED-ANOVA is robust even under smaller evaluation budgets.
D.4. Comparison with Existing Methods
In Section 5.4, we reported results for baselines computed with naive extensions of existing methods on the conditional-activation objective with the disjoint-domain setting (Equations 12 and 13). In this section, we present these baseline results for the conditional-activation objective with the overlapping-domain setting (Equations 12 and 14) and for the regime-dependent domain setting (Equations 15 and 16). The results are shown in Figures 9 and 10.
(c) Regime-dependent domains (Equation 15 and Equation 16)
First, for the conditional-activation objective, we observe the same trend as in the disjoint-domain setting (Equations 12 and 13). PED-ANOVA with filtering fails to distinguish between $x$ and $y$, assigning them nearly identical HPI values across all $\gamma^{\prime}$. This is clearly problematic, since the contributions of $x$ and $y$ to the objective should differ substantially. PED-ANOVA with imputation assigns $y$ an importance comparable to $c$ in the small-$\gamma^{\prime}$ regime, where $y$ should be inactive and thus cannot improve performance; it also exhibits unnatural instability for $\gamma^{\prime}>0.5$. f-ANOVA with filtering and MDI with filtering again yield nearly identical HPI values for all hyperparameters, and the remaining baselines also tend to assign almost the same importance to $x$ and $y$, failing to capture the true difference in their contributions to the objective.
Next, for the regime-dependent domain setting, PED-ANOVA with expansion also behaves undesirably: although $c$ determines whether a configuration falls into the top half, the importance of $c$ around $\gamma^{\prime}\approx 0.5$ becomes smaller than that of $x$ and $y$. This failure is caused by expansion, which collapses the regime-specific ranges of $x$ and $y$ into a unified domain, preventing accurate estimation. Moreover, for f-ANOVA, MDI, and SHAP with expansion, $y$ is assigned larger importance than $x$, even though $x$ should contribute more to the objective. This is because forcibly expanding the domain distorts the relationship between the variables and the objective across regimes, leading to misleading HPI estimates.
These results indicate that the naive extensions do not provide meaningful HPI estimates, not only for the conditional-activation objective under the disjoint-domain setting (as shown in Section 5.4.2), but also under the overlapping-domain setting and in the regime-dependent domain setting.
D.5. Ablation Study on the Effect of Regime-Level Weighting
In Figure 4, we showed the behavior of the $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ under the naive aggregation scheme (Equation 18) (bottom row of Figure 4). For reference, we plot $v_{\gamma,\mathrm{within}}^{(d)}$ used by condPED-ANOVA (which is normalized to obtain conditional local HPI) in Figure 11.
From this, we see that condPED-ANOVA applies appropriate regime-probability weighting, which calibrates the scale of $v_{\gamma,\mathrm{within}}^{(d)}$ for each variable and keeps it stable across regime changes without large discontinuities.
Appendix E Additional Experiments on Synthetic Problems
In Section 5, we considered the simplest synthetic problems. In this section, we study more complex synthetic settings.
E.1. Nested Conditional Activation
Figure 12. condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives on nested conditional activations (Equations 33 and 34). The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
E.1.1. Problem Setting
We next consider a synthetic problem with a nested conditional search space, where the active variable is selected through a two-level gating hierarchy. The objective function is defined as:
(33) $f(c_{0},c_{1},x,y,z)=\begin{cases}x,&\text{if }c_{0}<0.75\text{ and }c_{1}<0.5,\\ y,&\text{if }c_{0}<0.75\text{ and }c_{1}\geq 0.5,\\ z,&\text{if }c_{0}\geq 0.75,\end{cases}$
with the following variable domains:
(34) $\begin{aligned} c_{0}&\in[0,1],\\ c_{1}&\in[0,1],\end{aligned}\qquad\begin{aligned} x&\in[-7,-4],\\ y&\in[-3,0],\\ z&\in[2,5].\end{aligned}$
Our goal is to minimize $f$. When $c_{0}<0.75$, a second gating variable $c_{1}$ becomes active and selects between $x$ (if $c_{1}<0.5$) and $y$ (if $c_{1}\geq 0.5$); when $c_{0}\geq 0.75$, only the variable $z$ is active, and $c_{1}$, $x$, and $y$ are inactive.
E.1.2. Results and Discussion
The results are shown in Figure 12. We observe that $c_{0}$ attains the highest HPI around $\gamma^{\prime}\approx 0.75$, taking an importance close to one. This is reasonable, as $c_{0}$ induces the primary split of the search space. We also see that $c_{1}$ becomes important around $\gamma^{\prime}\approx 0.375$. Although its peak is lower than that of $c_{0}$ (around 0.75), this correctly reflects that, conditional on $c_{0}<0.75$, $c_{1}$ provides the next major bifurcation. Moreover, the importance shifts to the active variable within each regime: $x$ becomes large for $\gamma^{\prime}<0.375$, $y$ for $0.375<\gamma^{\prime}<0.75$, and $z$ for $\gamma^{\prime}>0.75$, with their HPI increasing as $\gamma^{\prime}$ moves away from the corresponding boundaries. These patterns are consistent with the nested conditional structure, indicating that condPED-ANOVA yields sensible importances even in nested conditional settings.
E.2. Combined Conditional Activation and Domain Shifts in a Three-Way Branching
E.2.1. Problem Setting
We next consider a synthetic problem with a three-way branching conditional search space, where the gating variable induces both conditional activation and regime-dependent domain shifts. The objective function is defined as:
(35) $f(c,x,y)=\begin{cases}x,&\text{if }c<1/3,\hskip 26.03424ptx\in[-7,-4],\\ x,&\text{if }1/3\leq c<2/3,\hskip 2.84526ptx\in[-3,0],\\ y,&\text{if }2/3\leq c,\hskip 26.74556pty\in[4,7],\end{cases}\ c\in[0,1].$
Our goal is to minimize $f$. Here, $x$ is always active when $c<2/3$, but its domain depends on the regime induced by $c$; when $c\geq 2/3$, $y$ becomes active and $x$ is inactive.
Figure 13. condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives on the combined conditional activation and domain shifts in a three-way branching setting (Equation 35). The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
E.2.2. Results and Discussion
The results are shown in Figure 13.
We observe that $c$ exhibits pronounced peaks around $\gamma^{\prime}\approx 0.33$ and $\gamma^{\prime}\approx 0.66$. This behavior is reasonable, since it is consistent with the role of $c$ in partitioning the objective-value range into three regimes. For $x$, the HPI is large in the regions $\gamma^{\prime}<0.33$ and $0.33<\gamma^{\prime}<0.66$, while it becomes nearly zero for $\gamma^{\prime}>0.66$. This pattern reflects that $x$ is influential when $c<1/3$ and when $1/3\leq c<2/3$. The near-zero HPI for $\gamma^{\prime}>0.66$ is also reasonable, as it corresponds to the regime where $x$ is inactive. Finally, the HPI of $y$ increases only in the $\gamma^{\prime}>0.66$ region and remains close to zero elsewhere. This is consistent with the fact that $y$ is active only in the regime $c\geq 2/3$.
These results suggest that condPED-ANOVA correctly computes HPI for branching hyperparameters such as $c$ that induce three or more regimes, as well as for hyperparameters such as $x$ whose activation and domain both vary across regimes.